OpenStudy (goldie1962):

Equation 1) p = kl^(1/2) Equation 2) T = 2π√L/g In equation 1, p = period and l = length. In equation 2, T = pendulum period, L = pendulum length, and g = gravity. This is the equation for the period of a pendulum. How does equation 1 relate to equation 2?

OpenStudy (loser66):

square both sides of both equations, you have \(p^2 = kl\\T^2 = 2\pi l/g\) if you solve for l from the first one and then replace to l on the second one, you have the relationship between p and T

OpenStudy (goldie1962):

So l = p^2 / k

OpenStudy (loser66):

yup

OpenStudy (goldie1962):

Thank you.

OpenStudy (loser66):

np